# Geodesic Trees and Exceptional Directions in FPP on Hyperbolic Groups

### v1

@@ Abstract.
We continue the study of the geometry of infinite geodesics in first passage percolation (FPP) on Gromov-hyperbolic groups $G$, initiated in [BT17] and developed further in [BM22]. It was shown in [BM22] that, given any fixed direction $\xi\in\partial G$, and under mild conditions on the passage time distribution, there exists almost surely a unique semi-infinite FPP geodesic from each $v\in G$ to $\xi$. Also, these geodesics coalesce to form a tree. Our main topic of study is the set of (random) exceptional directions for which uniqueness or coalescence fails. We study these directions in the context of two random geodesics trees: one formed by the union of all geodesics starting at a given base point, and the other formed by the union of all semi-infinite geodesics in a given direction $\xi\in\partial G$. We show that, under mild conditions, the set of exceptional directions almost surely has a strictly smaller Hausdorff dimension than the boundary, and hence has measure zero with respect to the Patterson-Sullivan measure. We also establish an upper bound on the maximum number of disjoint geodesics in the same direction. For groups that are not virtually free, we show that almost surely exceptional directions exist and are dense in ${\partial G}$. When the topological dimension of ${\partial G}$ is greater than one, we establish the existence of uncountably many exceptional directions. When the topological dimension of ${\partial G}$ is $n$, we prove the existence of directions $\xi$ with at least $(n+1)$ disjoint geodesics. Our results hinge on deep facts about hyperbolic groups, in particular on establishing the existence of random Cannon-Thurston maps for the geodesic trees. En route, we also establish facts about the structure of random bigeodesics that substantially strengthen prior results from [BT17]. The questions answered here have been considered before for FPP on euclidean lattices, but have only been satisfactorily answered for planar exactly solvable models of last passage percolation. Our results exhibit new features in higher dimensions.
@@ 1. Introduction
Over the last decade or two, there has been growing interest in understanding the interplay between the randomness of the metric and the underlying background geometry. FPP has been studied in different background geometries, typically on Cayley graphs of finitely generated groups (see e.g. [BT15, AG23, Gor24]). The study of FPP on Gromov hyperbolic groups was initiated in [BZ12, BT17] and was later continued in [BM22]. It was shown that while in some respects FPP on hyperbolic groups exhibit similar behavior to what is known or expected in the Euclidean (more specifically planar Euclidean) case, in other respects it shows sharply contrasting behavior. In this paper, we continue this investigation. We focus on the study of the structure of the geodesics. We answer several questions, analogues of which have been studied in the Euclidean case, but so far been satisfactorily answered only in planar exactly solvable models.
@@
For both forward and backward geodesic trees, of particular interest is the behavior of exceptional random directions, i.e.  random directions which show atypical behavior. For example, for the forward tree, recall that in any given direction $\xi$, there is almost surely a unique semi-infinite geodesic. However, since the number of directions is uncountable this allows for the existence of random directions where there might be multiple semi-infinite disjoint geodesics. Such directions are referred to as exceptional directions (a similar notion works for backward trees as well, precise definitions will be given later). Many natural questions arise: Are there exceptional directions? How many are there? How many disjoint geodesics can there be in an exceptional direction? These questions have been investigated in the Euclidean setting but the complete answer has so far been obtained only in the set-up of planar exponential last passage percolation (LPP) model and its scaling limit the directed landscape. We shall briefly recall the state of the literature in Section 1.2 below. For hyperbolic FPP, under mild conditions on the edge length distribution, we provide an almost complete answer to these questions (see Theorem 1.19 for a summary of the statements). By way of our proof, we also obtain results of independent interest about the structure of bigeodesics and geodesic trees which significantly strengthens previous results obtained in [BT17, BM22].
@@ 1.1. Basic set-up and statement of main results
  • $\partial G$–boundary of $G$.
  • $\nu$–Patterson-Sullivan measure on $\partial G$.
@@
We next recall some relevant results from [BM22]
@@
The following theorem says that almost surely all $\omega$-geodesic rays have a direction and that every $\xi\in{\partial G}$ is the direction of an $\omega$-geodesic ray..
@@ 1.2. Known results in Euclidean FPP and LPP
In this section we recall the relevant known results in Euclidean background geometry for those unfamiliar with Euclidean FPP. These results, specifically the ones known in planar exactly solvable LPP models, believed to be in the same universality class as planar Euclidean FPP, served as a motivation from the Euclidean planar setup for us to address the analogous issues in a hyperbolic context. The actual results, however, will not be used except to provide context in the rest of the paper. We start with describing the state of the art for FPP on Euclidean lattices.
@@
Euclidean FPP. The interest in infinite geodesics in FPP go back forty years. In the 1980s, the question, attributed to Furstenberg, whether or not a bi-infinite geodesic exists in Euclidean FPP was asked. This question remains open to date. The study of existence, uniqueness and coalescence of semi-infinite geodesics in given directions was taken up by Newman [New95] and his collaborators since the mid nineties and has received extensive attention since then. However, this question is intimately related to the properties of the limit shape for which we do not yet have an adequate understanding for FPP model on Euclidean lattices with generic edge length distributions. Therefore many of the existing results are conditional on various assumptions on the limit shape. In the planar case, some recent progress without such unverified assumptions have been made using the techniques of Busemann functions, [Hof05, Hof08, DH14, DH17, AH] being a representative sample of works in this direction. In particular, it is now known [DH14, DH17, AH] that, assuming differentiability of the boundary of limit shape, in every fixed direction uniqueness and coalescence of semi-infinite geodesics hold. In contrast to Theorem 1.11, it was shown in [DH14, Theorem 1.10(4)] that, under mild conditions, “backward clusters” are almost surely finite. In [AH], it was shown that there exists a family of random coalescing geodesics corresponding to linear functionals that are supporting to the limit shape. It is also known that almost surely there do not exist any bigeodesics in a fixed direction. However, this does not rule out existence of bigeodesics in random directions. Some recent results on coalescence under assumptions on the limit shape that are verifiable in certain cases were obtained in [DEP24]. We refer the interested reader to [ADH17, Chapter 4] for more details on the classically known results. Much of the progress in recent years in planar FPP is contained in [AHH22] and the survey paper [Ahl21]. Virtually nothing, however, is known in the higher dimensional setting.
@@
Exactly solvable models. As mentioned above, the question of existence, cardinality and multiplicity of exceptional directions are poorly understood in FPP, even in the planar setting. However, these questions are completely answered in the case of directed exponential LPP on $\mathbb{Z}^{2}$. In exponential LPP, one puts i.i.d. exponentially distributed weights on the vertices of $\mathbb{Z}^{2}$ and calculates the last passage time between two points by maximizing the sum of vertex weights over all up/right paths joining the two. Although last passage times form an anti-metric rather than a metric (and geodesics are defined as weight maximizing paths), it is believed that this model lies in the same KPZ universality class as planar FPP (under mild conditions). The following results are known in the case of exponential LPP (from universality considerations, one might believe the same results hold for planar FPP as well under some reasonable assumptions on the edge length distribution).
@@
In the hyperbolic context, the only result we are aware of is by Coupier, Flammant and Tran [CFT23] answering the N3G problem for the Hyperbolic Radial Spanning Tree, a random tree in the hyperbolic plane ${\mathbf{H}}^{2}$.
@@ 1.3. Novelty and structure of the paper
As described above, the existing literature focuses heavily on the planar context with the Euclidean background geometry, and the sharp results rely crucially on some form of exact solvability (the weakest form of the exact solvability being a product structure in the stationary version of the corresponding LPP processes). Our results, for the hyperbolic case, have a significantly broader scope: they utilize substantially different techniques and apply to a very large class of weight distribution and also to higher dimensional setting.
Theorem 1.19 (3) is indeed a close analog of the results described above in the Euclidean planar set-up, except that we are unable to completely solve the N3G problem. It is probably not reasonable to expect an exact analog of the N3G problem to go through in general without further assumptions. The existing solutions of the (planar Euclidean) N3G problem rely crucially on the integrability of the corresponding models (either via a connection to TASEP as in [Cou11] or exact computation of probability of three disjoint geodesics across a trapezium as in [Bus24]). We thus obtain a weaker but more general bound in Theorem 1.19 (6). In higher dimensions, we also establish behavior that contrasts with the planar exactly solvable cases (Theorem 1.19, (2), (4)): we show that the cardinality of the exceptional directions can be uncountable and the maximum multiplicity can be arbitrarily large depending on the topological dimension of ${\partial G}$.
@@ Acknowledgments
RB is partially supported by a SERB MATRICS grant from DST Govt. of India (MTR/2021/000093), and DAE project no. RTI4001 via ICTS. MM is partly supported by a grant from the TNQ foundation under the ”Numbers and Shapes” initiative, a DST JC Bose Fellowship, and by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500. Both authors were supported in part by an endowment of the Infosys Foundation. This project began at the International Centre for Theoretical Sciences (ICTS) during the program - Probabilistic Methods in Negative Curvature (Code: ICTS/PMNC2019/03). We gratefully acknowledge the support of ICTS.
@@ 2.1. Forward and backward trees with complete ends
Let $\omega\in{\mathcal{A}}_{\infty}$. Then, for any $\xi\in\partial G$ and any word geodesic $[1,\xi)$ defined by the sequence of vertices $1,v_{1},\cdots,v_{n}$, the only accumulation point of $[1,v_{n}]_{\omega}$ on $\partial G$ is $\xi$.

### v2

@@ Abstract.
We continue the study of the geometry of infinite geodesics in first passage percolation (FPP) on Gromov-hyperbolic groups $G$, initiated in [BT17] and developed further in [BM22]. It was shown in [BM22] that, given any fixed direction $\xi\in\partial G$, and under mild conditions on the passage time distribution, there exists almost surely a unique semi-infinite FPP geodesic from each $v\in G$ to $\xi$. Also, these geodesics coalesce to form a tree. Our main topic of study is the set of (random) exceptional directions for which uniqueness or coalescence fails. We study these directions in the context of two random geodesics trees: one formed by the union of all geodesics starting at a given base point, and the other formed by the union of all semi-infinite geodesics in a given direction $\xi\in\partial G$. We show that, under mild conditions, the set of exceptional directions almost surely has a strictly smaller Hausdorff dimension than the boundary, and hence has measure zero with respect to the Patterson-Sullivan measure. We also establish an upper bound on the maximum number of disjoint geodesics in the same direction. For groups that are not virtually free, we show that almost surely exceptional directions exist and are dense in ${\partial G}$. When the topological dimension of ${\partial G}$ is greater than one, we establish the existence of uncountably many exceptional directions. When the topological dimension of ${\partial G}$ is $n$, we prove the existence of directions $\xi$ with at least $(n+1)$ disjoint geodesics. Our results hinge on deep facts about hyperbolic groups, in particular on establishing the existence of random Cannon-Thurston maps for the geodesic trees. En route, we also establish facts about the structure of random bigeodesics that substantially strengthen prior results from [BT17].
@@ 1. Introduction
Over the last decade or two, there has been growing interest in understanding the interplay between the randomness of the metric and the underlying background geometry. FPP has been studied in different background geometries, typically on Cayley graphs of finitely generated groups (see e.g. [BT15, AG23, Gor24]). The study of FPP on Gromov hyperbolic groups was initiated in [BZ12, BT17] and was later continued in [BM22]. It was shown that while in some respects FPP on hyperbolic groups exhibit similar behavior to what is known or expected in the Euclidean (more specifically planar Euclidean) case, in other respects it shows sharply contrasting behavior. In this paper, we continue this investigation. We focus on the study of the structure of the geodesics. We answer a number of questions, analogs of which have been studied in the discrete Euclidean case, but so far been answered only in some planar exactly solvable models or some continuum models with additional rotational symmetries. As far as we are aware, our results are the first ones for FPP in a hyperbolic background geometry and are also the first ones in discrete models that are not exactly solvable. Our results also exhibit some qualitatively new features in the higher dimensional setting.
@@
For both forward and backward geodesic trees, of particular interest is the behavior of exceptional random directions, i.e.  random directions which show atypical behavior. For example, for the forward tree, recall that in any given direction $\xi$, there is almost surely a unique semi-infinite geodesic. However, since the number of directions is uncountable this allows for the existence of random directions where there might be multiple disjoint semi-infinite geodesics. Such directions are referred to as exceptional directions (a similar notion works for backward trees as well, precise definitions will be given later). Many natural questions arise: Are there exceptional directions? How many are there? How many disjoint geodesics can there be in an exceptional direction? These questions have been investigated in the Euclidean setting but the complete answer has so far been obtained only in the set-up of planar exponential last passage percolation (LPP) model and its scaling limit the directed landscape. We shall briefly recall the state of the literature in Section 1.2 below. For hyperbolic FPP, under mild conditions on the edge length distribution, we provide an almost complete answer to these questions (see Theorem 1.19 for a summary of the statements). By way of our proof, we also obtain results of independent interest about the structure of bigeodesics and geodesic trees which significantly strengthens previous results obtained in [BT17, BM22].
@@ 1.1. Basic set-up and statement of main results
  • $\partial G$– Gromov boundary of $G$, given by asymptote classes of geodesics in the Cayley graph.
  • $\nu$–Patterson-Sullivan measure on $\partial G$, given by a weak limit of uniform measure on $n$-spheres in the Cayley graph as $n\to\infty$.
@@
Notice that we have chosen not to formally define several standard objects in hyperbolic geometry, e.g., the Gromov boundary and the Patterson-Sullivan measure. We refer the reader to [BM22, Section 2.2] where relevant definitions and basic properties are are collected.
@@
We next recall some relevant results from [BM22].
@@
The following theorem says that almost surely all $\omega$-geodesic rays have a direction and that every $\xi\in{\partial G}$ is the direction of an $\omega$-geodesic ray.
@@
For the forward tree, under further assumptions on the group $G$, we can strengthen Theorem 1.19(2) further, and give a lower bound on the topological dimension of the set of exceptional directions; see Theorem 5.5.
@@ 1.2. Known results in Euclidean FPP and LPP
Some analogs of questions answered in Theorem 1.19 has been studied previously in various contexts, starting with FPP on Euclidean lattices, a notoriously hard model. As described below, there has been progress on this front, but sharper results have mostly been restricted to two related classes of models: planar exactly solvable models of last passage percolation, and certain models in the continuum (Euclidean $\mathbb{R}^{2}$ and $\mathbb{R}^{m}$) with additional rotational symmetries. In this section, we recall these results. These results, specifically the ones known in planar exactly solvable LPP models, believed to be in the same universality class as planar Euclidean FPP, served as a motivation from the Euclidean planar setup for us to address the analogous issues in a hyperbolic context. The actual results, however, will not be used except to provide context in the rest of the paper. We start with describing the state of the art for FPP on Euclidean lattices.
@@
FPP on Euclidean lattices. The interest in infinite geodesics in FPP go back forty years. In the 1980s, the question, attributed to Furstenberg, whether or not a bi-infinite geodesic exists in Euclidean FPP was asked. This question remains open to date. The study of existence, uniqueness and coalescence of semi-infinite geodesics in given directions was taken up by Newman [New95] and his collaborators since the mid nineties and has received extensive attention since then. However, this question is intimately related to the properties of the limit shape for which we do not yet have an adequate understanding for FPP model on Euclidean lattices with generic edge length distributions. Therefore many of the existing results are conditional on various assumptions on the limit shape. In the planar case, some recent progress without such unverified assumptions have been made using the techniques of Busemann functions, [Hof05, Hof08, DH14, DH17, AH] being a representative sample of works in this direction. In particular, it is now known [DH14, DH17, AH] that, assuming differentiability of the boundary of limit shape, in every fixed direction uniqueness and coalescence of semi-infinite geodesics hold. In contrast to Theorem 1.11, it was shown in [DH14, Theorem 1.10(4)] that, under mild conditions, “backward clusters” are almost surely finite. In [AH], it was shown that there exists a family of random coalescing geodesics corresponding to linear functionals that are supporting to the limit shape. It is also known that almost surely there do not exist any bigeodesics in a fixed direction. However, this does not rule out existence of bigeodesics in random directions. Some recent results on coalescence under assumptions on the limit shape that are verifiable in certain cases were obtained in [DEP24]. We refer the interested reader to [ADH17, Chapter 4] for more details on the classically known results. Much of the progress in recent years in planar FPP is contained in [AHH22] and the survey paper [Ahl21]. Virtually nothing, however, is known in the higher dimensional setting.
@@
Exactly solvable models. As mentioned above, the question of existence, cardinality and multiplicity of exceptional directions are poorly understood in lattice FPP, even in the planar setting. However, these questions are completely answered in the case of directed exponential LPP on $\mathbb{Z}^{2}$. In exponential LPP, one puts i.i.d. exponentially distributed weights on the vertices of $\mathbb{Z}^{2}$ and calculates the last passage time between two points by maximizing the sum of vertex weights over all up/right paths joining the two. Although last passage times form an anti-metric rather than a metric (and geodesics are defined as weight maximizing paths), it is believed that this model lies in the same KPZ universality class as planar FPP (under mild conditions). The following results are known in the case of exponential LPP (from universality considerations, one might believe the same results hold for planar FPP as well under some reasonable assumptions on the edge length distribution).
@@
Analogous results also exist for exactly solvable positive temperature polymer models, e.g. the log-Gamma polymer, see [JRA20, BFS24].
Models with additional continuous symmetries. As already mentioned, a key obstacle in understanding FPP on $\mathbb{Z}^{d}$ is the lack of control on the limit shape. To circumvent this, a number of models on (Euclidean) $\mathbb{R}^{d}$ have been considered with additional continuous rotational symmetry. This ensures that the limit shape is a Euclidean ball. One such model, based on a Poisson point process on $\mathbb{R}^{d}$, often referred to as the Howard-Newman model was introduced and studied in [HN97, HN01]. Using the rotational invariance of this model, the authors established an analog of Theorem 1.10 and showed that for the forward tree exceptional directions exist and are dense in $S^{d-1}$ [HN01, Theorem 1.10]. The question of multiplicity of exceptional directions was not considered in [HN01].
Another related result (but for a quite different model) in a non-lattice setting, and the only prior result we are aware of in a hyperbolic background geometry is by Coupier, Flammant and Tran [CFT23]. There, the authors consider a random tree constructed out of a Poisson process in the hyperbolic plane ${\mathbf{H}}^{2}$. This model is called the Hyperbolic Radial Spanning Tree. For this tree, they show that exceptional directions exist and are dense. They also solve the N3G problem in this context. We should point out, however, that this tree is not a geodesic tree constructed out of a first passage percolation model.
@@ 1.3. Novelty and structure of the paper
As pointed out above, existing results in the literature, mostly in the Euclidean background geometry, focus primarily on either exactly solvable models, or non-lattice models with additional continuous rotational symmetry. Such additional symmetries are typically not present in a Cayley graph. Our results appear to be the first in the lattice setting for a model which is not exactly solvable, as also the first in hyperbolic FPP. They apply to a large class of edge length distributions as well as in a higher dimensional setting. Existing literature on non-lattice models only deals with the case where the set of directions is a sphere $S^{d-1}$, whereas we are able to deal with more general boundaries ${\partial G}$ by appealing to deep facts about hyperbolic groups and the technology of Cannon-Thurston maps. Theorem 1.19 (3) is indeed a close analog of the results described above in the planar exactly solvable set-up, except that we are unable to completely solve the N3G problem. The existing solutions of the (planar Euclidean) N3G problem rely crucially on the integrability of the corresponding models (either via a connection to TASEP as in [Cou11] or an exact computation of probability of three disjoint geodesics across a trapezium as in [Bus24]). Instead we obtain a weaker but more general bound in Theorem 1.19 (6). We also establish features that appear to be qualitatively new: existence of exceptional directions with arbitrarily large multiplicity (Theorem 1.19,(4)) and an upper bound on the Hausdorff dimension of the set of exceptional directions (Theorem 1.19,(5)).
@@ Acknowledgments
The authors thank Ofer Busani for useful comments on a previous draft of this article. RB is partially supported by a SERB MATRICS grant from DST Govt. of India (MTR/2021/000093), and DAE project no. RTI4001 via ICTS. MM is partly supported by a grant from the TNQ foundation under the ”Numbers and Shapes” initiative, a DST JC Bose Fellowship, and by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500. Both authors were supported in part by an endowment of the Infosys Foundation. This project began at the International Centre for Theoretical Sciences (ICTS) during the program - Probabilistic Methods in Negative Curvature (Code: ICTS/PMNC2019/03). We gratefully acknowledge the support of ICTS.
@@ 2.1. Forward and backward trees with complete ends
Let $\omega\in{\mathcal{A}}_{\infty}$. Then, for any $\xi\in\partial G$ and any word geodesic $[1,\xi)$ defined by the sequence of vertices $1,v_{1},\cdots,v_{n},\ldots$, the only accumulation point of $[1,v_{n}]_{\omega}$ on $\partial G$ is $\xi$.
@@ 5.1. One and two-dimensional cases
5.2. Poincare duality groups and exceptional directions
The aim of this subsection is to strengthen Theorem 1.19 for hyperbolic Poincare duality groups. In [HN01] (see Remark following [HN01, Theorem 1.10]), the authors assert that in the Howard-Newman model with continuous rotational symmetry, exceptional directions in $\mathbb{R}^{n+1}$ must have Hausdorff dimension at least $n-1$. We were motivated, in part, by this assertion and provide below an analogous statement for hyperbolic Poincare duality groups (see [BM91] and [Mj12, Section 1.4] for relevant background). The following is the main theorem of this subsection.
Theorem 5.5.
Let $G$ be a hyperbolic $PD(n+1)$ group. Then, for a.e. $\omega$, the set of exceptional directions in the forward tree $F(1,\omega)$ has topological dimension at least $(n-1)$. Hence the Hausdorff dimension of the set of exceptional directions is at least $(n-1)$ almost surely. If, in addition, $G$ is the fundamental group of a closed negatively curved manifold, so that ${\partial G}=S^{n}$, then the set of exceptional directions in $F(1,\omega)$ has topological dimension exactly $(n-1)$ almost surely.
We start with a topological lemma.
Lemma 5.6.
Let $M$ be a closed orientable $n-$manifold, $A,B\subset M$ be closed sets, both with non-empty interior such that $A\cup B=M$. Let $C=A\cap B$. Further, assume that there exist interior points $a,b$ in the complements of $B,A$ respectively. Then $C$ has topological dimension at least $n-1$.
Proof.
Since $M$ is a manifold, $H_{n}(M,M-a,\mathbb{Z})=\mathbb{Z}=H_{n}(M,M-b,\mathbb{Z})$. Further, since $a,b$ are interior points of the closed sets $A,B$ respectively, $H_{n}(A,A-a,{\mathbb{Z}})={\mathbb{Z}}=H_{n}(B,B-b,{\mathbb{Z}})$. So, ${\mathbb{Z}}\subset H_{n}(A,C,{\mathbb{Z}})$ and ${\mathbb{Z}}\subset H_{n}(B,C,{\mathbb{Z}})$.
We argue by contradiction. Suppose that $C$ has topological dimension at most $(n-2)$. We now use the fact [Wal81] that if the topological dimension of a set is finite, then the topological dimension equals the Cech (co)homological dimension over $\mathbb{Z}$, so that $H_{n-1}(C,{\mathbb{Z}})=0$. Then the homology exact sequence of $(A,C)$ forces $H_{n}(A,C,{\mathbb{Z}})=H_{n}(A,\mathbb{Z})$. Similarly, for the pair $(B,C)$.
Now, applying the Mayer-Vietoris for $M,A,B$, we find that since $C=A\cap B$ has topological dimension at most $(n-2)$, $H_{n}(M)=H_{n}(A)\oplus H_{n}(B)$ and hence $\mathbb{Z}\oplus\mathbb{Z}<H_{n}(M)$, a contradiction. ∎
Note that the above argument used only the algebraic facts that $H_{n}(M,M-a,\mathbb{Z})=\mathbb{Z}=H_{n}(M,M-b,\mathbb{Z})$, and $H_{n}(M)=\mathbb{Z}$. These are exactly the defining conditions of a homology manifold. Recall that $M$ is said to be an $n-$dimensional homology manifold if
  1. (1)
    $M$ is a locally compact Hausdorff space with finite homological dimension over $\mathbb{Z}$.
  2. (2)
    For all $m\in M$, $H_{n}(M,M\setminus\{m\},\mathbb{Z})=\mathbb{Z}$ and $H_{i}(M,M\setminus\{m\},\mathbb{Z})=0$ for $i\neq n$.
  3. (3)
    $H_{n}(M,\mathbb{Z})=\mathbb{Z}$.
Thus, the proof of Lemma 5.6 goes through for $M$ a homology manifold verbatim. We record this:
Corollary 5.7.
Let $M$ be a homology manifold, and $A,B\subset M$ be closed sets, both with non-empty interior such that $A\cup B=M$. Let $C=A\cap B$. Further, assume that there exist interior points $a,b$ in the complements of $B,A$ respectively. Then $C$ has topological dimension at least $n-1$.
We now recall the following theorem of Bestvina-Mess (see [Mj12, Section 1.4] for a quick summary of the background material).
Theorem 5.8.
[BM91] Let $G$ be a hyperbolic $PD(n+1)$ group. Then $\partial G$ is a homology manifold of dimension $n$.
Proof of Theorem 5.5.
Let $G$ is a hyperbolic $PD(n+1)$ group and $\Gamma$ a Cayley graph with respect to a finite symmetric generating set. Recall the event $\mathcal{C}_{\varepsilon,r}$ from Lemma 4.5. By Lemma 4.5 and the Borel-Cantelli Lemma, lim $\liminf\mathcal{C}_{\varepsilon,r}$ is an almost sure event. For $\omega\in\liminf\mathcal{C}_{\varepsilon,r}$, let $r_{0}(\omega)$ be such that $\mathcal{C}_{\varepsilon,r}$ holds for all $r\geq r_{0}$. Now, for $r\geq r_{0}$ and $v\in\partial B_{r}(1)$, define the $\omega-$shadow $\mho(v,\omega)$ of $v$ to be the collection of $\xi\in{\partial G}$ such that there exists some $[1,\xi)_{\omega}$ satisfying the property that $v$ is the first point of intersection of $[1,\xi)_{\omega}$ with $\partial B_{r}(1)$. Observe that $\mho(v,\omega)$ is a closed subset of ${\partial G}$. Indeed, suppose $\xi_{n}\in\mho(v,\omega)$ such that $\xi_{n}\to\xi$. Since the restriction of $[1,\xi_{n})_{\omega}$ from $1$ to $v$ is contained in $B_{r}(1)$ for each $\xi_{n}$ it follows by a diagonal argument that there exist an infinite path $\{1=u_{0},u_{1},\ldots,\}$ from $1$ that is a subsequential limit of $[1,\xi_{n})_{\omega}$ that intersects $\partial B_{r}(1)$ first at $v$. By definition, each initial segment of this path is an $\omega$-geodesic and hence this path itself is also an $\omega$-geodesic ray. Since $\xi_{n}\to\xi$ it follows that $u_{i}\to\xi$ and therefore there exists a $[1,\xi)_{\omega}$ which intersects $\partial B_{r}(1)$ first at $v$. Therefore $\xi\in\mho(v,\omega)$.
On the event $C_{\varepsilon,r}$ it follows that if $\xi\in\mho(v,\omega)$, then $v\in B_{\varepsilon r}(w)$ where $w=[1,\xi)\cap\partial B_{r}(1)$. Therefore $\mho(v,\omega)$ is contained in the deterministic $\varepsilon r$-shadow $\mathcal{S}(v,\varepsilon r)$ of $v$. Notice that by restricting $\omega$ to a further almost sure subset if necessary, we can assume by Theorem 2.11 that ${\partial G}$ is a finite union of closed sets $\cup_{u\in\partial B_{r}(1)}\mho(v,\omega)$. Therefore, there exists $v\in\partial B_{r}(1)$ such that $\mho(v,\omega)$ has a non-empty interior (by the Baire category theorem, for instance). Also, notice that, by choosing $\varepsilon$ sufficiently small there exists $w\in\partial B_{r}(1)$ such that the deterministic shadows $\mathcal{S}(v,\varepsilon r)$ and $\mathcal{S}(w,\varepsilon r)$ are disjoint. Indeed, given $k\in{\mathbb{N}}$, if the Gromov inner product $\langle v,w\rangle_{1}\leq k$, then the deterministic shadows $\mathcal{S}(v,D^{\prime})$ and $\mathcal{S}(w,D^{\prime})$ are disjoint for $v,w\in\partial B_{r^{\prime}}(1)$ for $r^{\prime}\geq 2k+D+\delta$. Set $D=\varepsilon r$, and $r^{\prime}=r$. Then for $\varepsilon\in(0,1)$, one can choose $r$ large enough depending on $\varepsilon$ such that $r^{\prime}\geq 2k+D+\delta$. Therefore, for large enough $r$, there exists $v,w\in\partial B_{r}(1)$ such that $\mathcal{S}(v,\varepsilon r)$ and $\mathcal{S}(w,\varepsilon r)$ are disjoint, as claimed. Now, set $A=\mho(v,\omega)$, and let $B=\cup_{u\neq v,u\in\partial B_{n}(1)}\mho(u,\omega)$. Notice that $A$ and $B$ are closed subsets of ${\partial G}$ with non-empty interiors ($A$ has non-empty interior by choice, and $B$ has non-empty interior since it contains $\mathcal{S}(w,\varepsilon r)$). Also $A\cup B={\partial G}$.
Then, by Corollary 5.7 and Theorem 5.8, $A\cap B$ has topological dimension at least $n-1$. Observe now that $a\in A\cap B$ implies $a$ is an exceptional direction in $F(1,\omega)$. Indeed, any $a\in A\cap B$ belongs to $\mho(u,\omega)\cap\mho(v,\omega)$ for some $u$. Therefore there exist $[1,a)_{\omega}^{1}$ and $[1,a)^{2}_{\omega}$ whose first intersection points with $\partial B_{r}(1)$ are distinct. Thus, $a$ is an exceptional direction of $F(1,\omega)$. Therefore, the set of exceptional directions has topological dimension at least $n-1$.
To conclude the proof of Theorem 5.5, it suffices to show that the topological dimension of exceptional points cannot equal $n$ if ${\partial G}=S^{n}$. This follows from [HW41, Theorem IV 3] which asserts that if the topological dimension of a subspace of $\mathbb{R}^{n}$ is $n$, then it has non-empty interior. The same holds for $S^{n}$ as the latter is the one-point compactification of $\mathbb{R}^{n}$. This is not possible by Theorem 4.1 and Ahlfors regularity of ${\partial G}$ which implies that every open set in ${\partial G}$ must have positive Patterson-Sullivan measure. ∎
We conclude this subsection with the following general question (without extra hypotheses on $G$).
Question 5.9.
If $G$ is hyperbolic with ${\partial G}$ of topological dimension $n$, do the set of exceptional directions have topological dimension $(n-1)$?